OpenStudy (anonymous):

Help. Input the equation of the given line in standard form. The line with m = and the point (-2, -3).

OpenStudy (anonymous):

The eq of line in slope-point form is given as \[y-y_1= m(x-x_1)\] where \[x_1= -2 \] & \[y_1=-3\] Substituting these values in the eq of line we find \[y-(-3)= m[x-(-2)]\] \[y+3= m(x+2)\] or \[y= m(x+2) -3\] is the required eq of the line

OpenStudy (anonymous):

@DjNck